OpenStudy (anonymous):

A line passes through the points (-6, 3) and (x2, y2), and has a slope of 5/10. If point (x2, y2) is located in quadrant I, find x2. 4 -12 -7 -5

OpenStudy (anonymous):

any ideas

OpenStudy (anonymous):

theres one piece of info missing

OpenStudy (anonymous):

2y - 6 = x+6 x2 = 2y2 -12

OpenStudy (anonymous):

this is all the problem says

OpenStudy (anonymous):

u need the value of y2

OpenStudy (anonymous):

the answer was 4 i went ahead and guessed

OpenStudy (anonymous):

guessing wont do u good...in the long term

OpenStudy (anonymous):

ive been n childrens hospital. I have missed all the explaining and my teacher exspects me to know it all. and its harder then she thinks.

OpenStudy (anonymous):

no prob..dont worry...were to help

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

its obvious it had to be 4 lols , its the only positive answer, and for the point to be in the first quadrant then x must be positive

OpenStudy (anonymous):

oh that way..lol

OpenStudy (anonymous):

lol. How can they ask such questions? :D Neways good luck jessmaynard!